What is the pressure exerted by a .500mol sample of Nitrogen in a 10.0 Liter container at 293K
1
Expert's answer
2019-10-15T07:45:06-0400

Solution.

The equation of ideal gas:

"p \times V = n \times R \times T"

"p = \frac{n \times R \times T}{V}"

p = 121741.5 Pa

Answer:

p = 121741.5 Pa